Presented below are data relating to labor for Verde Appliance Repair Shop.

Repair-technicians' wages $110,000
Fringe benefits 26,400
Overhead 48,400


The desired profit margin per hour is $28. The material loading charge is 60% of invoice cost. Verde estimates that 4,400 labor hours will be worked next year. If Verde repairs a dishwasher that takes 1.5 hours to repair and uses parts of $80, compute the bill for the job.

Total cost $

Presented below are data relating to labor for Verde Appliance Repair Shop.

Repair-technicians' wages $110,400
Fringe benefits 23,000
Overhead 50,600


The desired profit margin per hour is $20. The material loading charge is 60% of invoice cost. Verde estimates that 4,600 labor hours will be worked next year. If Verde repairs a dishwasher that takes 1.3 hours to repair and uses parts of $60, compute the bill for the job.

Total cost

$enter the total cost in dollars
Answer #1
1
Budgeted cost per labor hour 42.00 =(110000+26400+48400)/4400
Add: Profit margin 28.00
Labor rate 70.00
Labor charges 105.00 =1.5*70
Materials parts 80.00
Material loading charges 48.00 =80*60%
Total cost 233.00
2
Budgeted cost per labor hour 40.00 =(110400+23000+50600)/4600
Add: Profit margin 20.00
Labor rate 60.00
Labor charges 78.00 =1.3*60
Materials parts 60.00
Material loading charges 36.00 =60*60%
Total cost 174.00